OpenStudy (spareb665):

Express ln√ab/c^3 as a sums and difference of logarithms.

OpenStudy (spareb665):



OpenStudy (anonymous):

\[ \ln\sqrt{\frac{ab}{c^3}} =\ln\frac{\sqrt{a}\sqrt{b}}{\sqrt{c^3}}\] Does it help?

OpenStudy (spareb665):

i just needed help not the answer :(

OpenStudy (anonymous):

I just showed you one step not the whole solution.

OpenStudy (spareb665):

I'm trying to understand how to solve these problems.

OpenStudy (anonymous):

Ok. In theese problem it is important to know that logarithm of a product is sum of logaritms. \[ \ln(xy) = \ln x + \ln y\]

OpenStudy (spareb665):

in(

OpenStudy (anonymous):

You can write \( \sqrt{abc} \) as \(\sqrt{a}\sqrt{b}\sqrt{c}\). Then you have a product of three terms.

OpenStudy (spareb665):

ok but isnt taht the same thing as abc square rooted?

OpenStudy (anonymous):

Yes, it is. Now you can express the log of a product as a sum of logs.

OpenStudy (anonymous):

Remember \[\ln(xy) = \ln x + \ln y\] \[ \ln{\left(\frac{x}{y}\right)} = \ln x - \ln y\] \[ \ln x^a = a \ln x \] These three formulas are what you have to know to solve this kind of problem.